# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v3

@@ 4.2.1 TF-IDF
TF-IDF (Term Frequency - Inverse Document Frequency) is a well-known statistical formula that calculates the importance of words relative to a corpus. A few works investigated the extraction of topological features from TF-IDF representations as part of the pipeline illustrated in Fig. 4. For instance, SIFT, a persistent homology-based model with TF-IDF, is developed to differentiate between child and adolescent writings Zhu (2013). This model represented the TF-IDF features as a time series, and then extracted topological features to enhance text classification. Several other researchers applied this model to other classification tasks, such as deepfake text detection Løvlie (2023), presidential election speech attribution Huang (2022), and movie genre classification Doshi and Zadrozny (2018); Shin (2019).
Insight (TFIDF). Topological features extracted from TF-IDF have only been applied to the text classification task. It is possible that topological features extracted from TF-IDF are not rich enough to perform the other six tasks.
@@ 4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Unlike TF-IDF, topological features extracted from word2vec have been applied to several tasks, including structure & visualization, where Haghighatkhah et al. (2022) creates story trees to trace story lines. Next, we observe applications in sentiment & semantic analysis task, where TDA is applied to novel problems such as the creation of a topological search engine Cornell (2020), measuring distance between the literary style of Spanish poets Paluzo Hidalgo et al. (2019), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), derivation of the correlation between sentence vectors and their semantics Sun and Nelson (2023), and comparing text and speech embeddings Yessenbayev and Kozhirbayev (2022, 2024). Thirdly, for the classification task, researchers detect fraudulent papers Tymochko et al. (2021), and topological loops in logical statements Tymochko et al. (2020).
Additionally, we observe applications in the health, social, and scholarly analysis task - disease prediction from epidemic curves Petri and Leitao (2020), topic modeling tasks Holmes (2020); Wright and Zheng (2020). Finally, for the model interpretation and analysis task, Feng et al. (2024) uses both topological and geometrical features to investigate the strength of LLM-enhanced data augmentation.
Insight (Word2Vec). Topological features extracted from Word2vec are considered rich, such that researchers have attempted six out of seven tasks. The only tasks that have not been attempted are speech and music processing, since different embedding is needed for audio data. Finally, we observe that the sentiment and semantic analysis task is the most popular task for word2vec.
@@
GloVe or Global Vectors for Word Representation is another technique for numerically representing texts as embeddings. The tasks we observe that are attempted using topological features extracted from GloVe embeddings are - classification, where subtasks include novel author attribution Gholizadeh et al. (2018) , deepfake text detection Løvlie (2023); sentiment and semantic analysis Gholizadeh et al. (2020); Tymochko et al. (2020); Zadrozny (2021b); model interpretation and analysis, where Haim Meirom and Bobrowski (2022); Michel et al. (2017) compare text representations and embeddings, Spannaus et al. (2024) explains model performance, and Zadrozny (2021a) tests the manifestation of intelligence and understanding in models; health, social, and scholarly analysis Byers (2021); Novak (2019); and clustering and topic modeling Deng and Duzhin (2022)
Insight (GloVe). Topological features extracted from GloVe are considered rich, such that researchers have attempted all tasks, except (1) speech and music processing and (2) structure and visualization. The most popular application is model interpretation and analysis, suggesting that glove embeddings can be sufficiently probed using TDA techniques to excavate explanations for performance.
@@
FastText embeddings build on the Word2Vec approach by incorporating subword information, improving the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). This type of embedding is not a popular feature researchers use to enhance topological features as only two tasks are attempted - sentiment and semantic analysis Jakubowski et al. (2020); Triki (2021); Shehu (2024), specifically for word sense induction & disambiguation; and text classification, where Tymochko et al. (2021) detects fraudulent papers.
Insight (FastText). FastText is not popularly adopted for topological applications, with only a few researchers applying it to two tasks - text classification and sentiment & semantic analysis.
@@ 4.2.3 Pre-trained Contextual Embeddings
Researchers evaluated the strength of the TDA features extracted from Transformer-based Vaswani et al. (2017) embeddings. Using the idea of self-attention, the neural network can encode more semantic and syntactic features than previous embeddings which should allow for richer TDA features to be extracted. To incorporate TDA features for various tasks, several researchers have investigated the efficacy of using other outputs of encoder and decoder models - CLS Embedding output, hidden weights, and attention weights to extract high-quality additional features.
CLS Embedding Output. Researchers have applied these features on text classification tasks, specifically for deepfake text detection Tulchinskii et al. (2024); Kushnareva et al. (2024); Wei et al. (2025); Guilinger et al. (2025), fake news detection Lavery et al. (2024), and TEDtalk public speaking ratings classification Das et al. (2021).
@@
Additionally, we observe applications to the topic modeling task Byrne et al. (2022), and the model interpretation and analysis task Gourgoulias et al. (2024); Proskura and Zaytsev (2024). Gourgoulias et al. (2024) probes LLMs to estimate class separability of text datasets, and Proskura and Zaytsev (2024) uses topological information from encoder models to select the best models to create an ensemble. Finally, Rathore et al. (2023) performs the structure and visualization task by visualizing the training process of transformer-based models.
@@
Insight (Transformer-CLS). CLS embedding is applied to four tasks - text classification (the most popular), topic modeling, model interpretation & analysis, and structure & visualization. The most interesting subtasks are probing LLMs to estimate class separability of text datasets, and visualizing the training process of transformer-based models.
@@
Hidden Weights. Classification tasks include deepfake text detection Uchendu et al. (2024), and LLM hallucination detection Bazarova et al. (2025). Next, researchers explore a combination of sentiment & semantic analysis and structure & visualization tasks by using Mapper to visualize polysemous words in the hidden representations of the BERT transformer model Garcia (2022). Similarly, Alexander and Wang (2023) combines the health analysis and visualization tasks to visualize GPT-3’s embeddings of hate speech, misinformation, and psychiatric disorder texts with Mapper. Also, Ruppik et al. (2024) performs the clustering and topic modeling task through dialogue term extraction. Finally, model interpretation and analysis is the most popular task, where Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to prune redundant layers in LLMs; Balderas et al. (2025) proposes Persistent BERT Compression and Explainability (PBCE) to compress BERT by pruning redundant layers; Sun and Nelson (2023) probes the correlation between sentence vectors and their semantics; Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT; García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification; and Kudriashov et al. (2024) probes BERT’s hidden weights on new grammatical features, known as polypersonality.
@@
Insight (Transformer-Hidden). Researchers have applied TDA features extracted from Hidden weights in to all tasks, except health, social, & scholarly analysis and speed & music processing tasks. From the various novel model interpretation and analysis applications, it seems that TDA features extracted from the Hidden weights are able to make these black-box models more transparent.
@@
Attention Weights. BERT’s attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted for text classification tasks such as deepfake text detection Kushnareva et al. (2021); Uchendu et al. (2024), robustness evaluation of TDA features Perez and Reinauer (2022), authorship attribution of Japanese texts Sakurai et al. (2025), out-of-distribution detection (OOD) Pollano et al. (2024); Perez and Reinauer (2022), and vulnerability detection in code Snopov and Golubinskiy (2024). Next, this framework is applied to the sentiment and semantic analysis task, specifically on human linguistic competence (i.e., grammatical acceptability judgment)  Cherniavskii et al. (2022); Proskurina et al. (2023); Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), and document coherence Jain et al. (2024). Similarly, with the same framework, a speech and music processing application Tulchinskii et al. (2023). Finally, researchers attempt the model interpretation and analysis task, where Kostenok et al. (2023) uses the topological features extracted from the attention weights to estimate uncertainty in the encoder models, and Proskura and Zaytsev (2024) performs dynamic weighting for creating ensemble models.
@@
Insight (Transformer-Attention). Given the wealth of information contained in the attention weights, researchers have applied topological features extracted from these weights to several tasks. The most interesting applications are observed on the sentiment and semantic tasks - grammatical acceptability judgment and document coherence analysis. Finally, we observe that for all the Transformer-based weights, the most explored application is deepfake text detection.
@@
ELMo Embeddings. ELMo embeddings are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) performs text classification by applying these embeddings for detecting fraudulent papers by examining their titles and abstracts. This is done in comparison of other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
@@
Insight (ELMo). ELMo embeddings is applied to one task - text classification by one researcher in comparison to other embeddings.
@@ 4.2.4 Symbolic Representations
Symbolic representations in the context of AI and cognitive science refer to the use of symbols such as letters, numbers, tokens, or abstract entities to represent concepts, objects, relationships, and rules within a system. These symbols can be manipulated according to predefined rules to perform reasoning, problem-solving, and decision-making. Symbolic representation contrasts with sub-symbolic representations, such as neural network-based embeddings, which do not explicitly use symbols or rules. This section then discusses the creation of symbolic representations by using principles of letter coding (PLC) and principles of speech sound coding (PSSC) which are then used to extract topological features.
PLC refers to rules and methods used to encode letters that fuel various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters, such as for Ukrainian Yurchuk and Gurnik (2023); Kovaliuk et al. (2024). These applications fall under the text classification task for PLC and speech & music processing task.
Insight (Symbolic). Symbolic representations are a novel technique for numerically representing text. The only applications of this technique are on Uranian tongue twisters, both for the classification and speech & processing tasks. This suggests that such a technique could be applied to processing low-resource languages, where numerical representation may be insufficient.
@@ 4.2.5 Multi-Modal Representations
TDA features have also been extracted from other representations of NLP-related features, including multimedia data such as audio and video. In this section, the only task performed by researchers is speech & music processing, where applications include - studying human vowels and infant vocalizations Bonafos et al. (2023, 2024), emotion recognition from audio speech Gonzalez-Diaz et al. (2019) and videos Paluzo-Hidalgo et al. (2022), depression detection from audio clips Tlachac et al. (2020), recognizing voiced and voiceless consonants in speech Zhu et al. (2024), music classification Bergomi (2015); Sassone et al. (2022), and assessing the adversarial robustness of image-text multi-modal models by measuring topological consistency Vu et al. (2025).
Insight (Multi-Modal). Multi-modal representations are another novel technique for numerically representing. The only task performed by researchers using this embedding is the speech & music processing task. This is because of the nature of the data being audio. Finally, the most interesting applications are the emotion recognition from audio speech, and depression detection from audio clips. All researchers show that topological features extracted from such embeddings experience high gains.

### v4

@@ 5.3.1 TF-IDF
TF-IDF (Term Frequency - Inverse Document Frequency) is a well-known statistical formula that calculates the importance of words relative to a corpus. A few works investigated the extraction of topological features from TF-IDF representations as part of the pipeline illustrated in Fig. 6. For instance, SIFT, a persistent homology-based model with TF-IDF, is developed to differentiate between child and adolescent writings (Zhu, 2013). This model represented the TF-IDF features as a time series, and then extracted topological features to enhance text classification. Several other researchers applied this model to other classification task, such as deepfake text detection (Løvlie, 2023), presidential election speech attribution (Huang, 2022), distinguishing between languages by averaging the persistence landscapes Sovdat (2016), age group categorization of lonely people (Effah, 2017), and movie genre classification (Doshi & Zadrozny, 2018; Shin, 2019). Additionally, Elyasi & Moghadam (2019) compares the two popular TDA approaches - Persistent Homology and Mapper to classify Persian poems. Also, using Mapper for the structure & visualization tasks, Maadarani & Hajra (2020) explains linguistic phenomena in poetry writing styles, and van Veen (2020) interprets NLP model behavior. Lastly, we observe applications in the clustering & topic modeling task - keyphrase extraction (Guan et al., 2016), text summarization (Kumar & Sarkar, 2022), and twitter topic detection (Torres-Tramón et al., 2015); sentiment & semantic analysis task - legal entailment (Savle et al., 2019), and sentiment analysis of movie reviews (Michel et al., 2017).
Insight (TFIDF). Topological features extracted from TF-IDF have been applied to four out of the seven tasks - classification, clustering & topic modeling, sentiment & semantic analysis, and structure & visualization. This suggests that while there have been an exponential improvement in contextual and non-contextual word embeddings for numerically representing texts, TF-IDF can still extract decent features from text that can be further leveraged by TDA.
@@ 5.3.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations (Mikolov et al., 2013). Thus, we observe applications in the structure & visualization task, where Haghighatkhah et al. (2022) creates story trees to trace story lines. Next, we observe applications in sentiment & semantic analysis task, where TDA is applied to novel problems such as the creation of a topological search engine using Mapper (Cornell, 2020), measuring distance between the literary style of Spanish poets - Francisco de Quevedo, Luis de Góngora, and Lope de Vega (Paluzo Hidalgo et al., 2019), detecting narrative shifts in media discourse (Bailey & Heiligman, 2025), analysis of contradictions within texts (Wu et al., 2022), and word sense induction and disambiguation (Rawson et al., 2022; Temčinas, 2018). Thirdly, for the classification task, researchers detect fraudulent papers (Tymochko et al., 2021), and topological loops in logical statements (Tymochko et al., 2020).
Furthermore, we observe applications in the health, social, and scholarly analysis task - disease prediction from epidemic curves (Petri & Leitao, 2020), topic modeling tasks (Holmes, 2020; Wright & Zheng, 2020). Finally, for the model interpretation and analysis task, Feng et al. (2024) uses both topological and geometrical features to investigate the strength of LLM-enhanced data augmentation, Sun & Nelson (2023) derives the correlation between sentence vectors and their semantics, and Yessenbayev & Kozhirbayev (2022; 2024) compares the semantic alignment of text and speech embeddings.
Insight (Word2Vec). Topological features extracted from Word2vec are considered rich, such that researchers have attempted six out of seven tasks. The only task that have not been attempted are speech and music processing, since different embedding is needed for audio data. This suggests that word2vec embeddings capture enough linguistic features that when augmented with topological features can improve baseline performance. Finally, we observe that the sentiment and semantic analysis task is the most popular task for word2vec.
@@
GloVe or Global Vectors for Word Representation is another technique for numerically representing texts as embeddings. Topological features extracted from GloVe embeddings have been applied to the following tasks - classification task, which include author attribution of novel authors (Gholizadeh et al., 2018), fake news detection (Deng & Duzhin, 2022), and deepfake text detection (Løvlie, 2023); sentiment & semantic analysis task, which include document categorization (Gholizadeh et al., 2020), and capturing circles in circular arguments (Tymochko et al., 2020; Zadrozny, 2021b); model interpretation and analysis, where Haim Meirom & Bobrowski (2022); Michel et al. (2017) compare text representations & embeddings, Spannaus et al. (2024) explains model performance, and Zadrozny (2021a) tests the manifestation of intelligence and understanding in models; and health, social, and scholarly analysis task - social anxiety detection (Byers, 2021), and keywords extraction of scholarly documents (Novak, 2019).
Insight (GloVe). Topological features extracted from GloVe are considered rich, such that researchers have attempted all tasks, except (1) speech & music processing, (2) structure & visualization, and (3) clustering & topic modeling. The most popular application is in the model interpretation & analysis task, suggesting that glove embeddings can be sufficiently probed using TDA techniques to excavate interpretations of model performance.
@@
FastText embeddings build on the Word2Vec approach by incorporating subword information, improving the representation of rare words, and allowing for embedding out-of-vocabulary words (Bojanowski et al., 2017). This type of embedding is not a popular feature extractor researchers employ to enhance topological features as only two tasks are attempted - sentiment & semantic analysis task, which include polysemy word classification (Jakubowski et al., 2020; Triki, 2021; Shehu, 2024), and word sense induction & disambiguation (Jakubowski et al., 2020); and text classification, where Tymochko et al. (2021) detects fraudulent papers.
Insight (FastText). FastText is not widely adopted for topological applications, with only a few researchers applying it to two tasks - text classification and sentiment & semantic analysis. Sentiment & Semantic analysis is the more popular application involving polysemy word classification and word sense induction & disambiguation.
@@ 5.3.3 Pre-trained Contextual Embeddings
Researchers have evaluated the strength of the TDA features extracted from Transformer-based (Vaswani et al., 2017) embeddings. Using the idea of self-attention, the neural network can encode more semantic and syntactic features than previous embeddings which should allow for richer TDA features to be extracted. To incorporate TDA features for various tasks, several researchers have investigated the efficacy of using other outputs of encoder and decoder Transformer models - CLS Embedding output, hidden weights, and attention weights to extract high-quality additional features.
@@
CLS Embedding Output. Researchers have applied these features on text classification task, specifically for deepfake text detection (Tulchinskii et al., 2024; Kushnareva et al., 2024; Wei et al., 2025; Guilinger et al., 2025), fake news detection (Lavery et al., 2024), and TEDtalk public speaking ratings classification (Das et al., 2021). Additionally, we observe applications to the topic modeling task (Byrne et al., 2022), and the model interpretation & analysis task (Gourgoulias et al., 2024; Proskura & Zaytsev, 2024). Gourgoulias et al. (2024) probes LLMs to estimate class separability of text datasets, and Proskura & Zaytsev (2024) uses topological information from encoder models to select the best models to create an ensemble. Furthermore, Arun et al. (2025) employs TDA for the structure & semantic task, detecting controversial vs. non-controversial political discourse by capturing the shifts in discourse for controversial data. Finally, Rathore et al. (2023) performs the structure & visualization task in combination with the model interpretation & analysis task by visualizing the training process of transformer-based models.
@@
Insight (Transformer-CLS). CLS embedding is applied to five tasks - text classification (the most popular), topic modeling, model interpretation & analysis, sentiment & semantic analysis, and structure & visualization. The most interesting subtasks are probing LLMs to estimate class separability of text datasets and visualizing the training process of transformer-based models.
@@
Hidden Weights. Classification task include deepfake text detection (Uchendu et al., 2024), and LLM hallucination detection (Bazarova et al., 2025). Next, Garcia (2022) explore a combination of the sentiment & semantic analysis and structure & visualization tasks by using Mapper to visualize polysemous words in the hidden representations of the BERT transformer model. In addition, Bensalem et al. (2025), performs a sentiment analysis task by using the sentiment scores of original vs. translated texts extracted from a Transformer-based model, representing these scores in a time series form and using zigzag persistent homology to detect sentiment shift in translated texts. Similarly, Alexander & Wang (2023) combines the health analysis and visualization tasks to visualize GPT-3’s embeddings of hate speech, misinformation, and psychiatric disorder texts with Mapper. Also, Ruppik et al. (2024) performs the clustering & topic modeling task through dialogue term extraction. Finally, we observe that the model interpretation & analysis task is the most popular task, where Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to prune redundant layers in LLMs; Balderas et al. (2025) proposes Persistent BERT Compression and Explainability (PBCE) to compress BERT by pruning redundant layers; Sun & Nelson (2023) probes the correlation between sentence vectors and their semantics; Chauhan & Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT; Fay et al. (2025) investigates the differences in the topological structure of the latent space of adversarial vs. non-adversarial texts in LLMs; García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification; Fitz (2023) investigates the topological structure of the brain of ChatGPT concerning its notion of fairness; and Kudriashov et al. (2024) probes BERT’s hidden weights on new grammatical features, known as polypersonality.
@@
Insight (Transformer-Hidden). Researchers have applied TDA features extracted from Hidden weights in to all tasks, except speech & music processing tasks. The most popular task attempted is the model interpretation & analysis task, suggesting that TDA features extracted from the Hidden weights are able to make these black-box models more transparent.
@@
Attention Weights. Attention weights extracted from BERT and its variants (i.e., RoBERTa) have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted for the text classification task such as deepfake text detection (Kushnareva et al., 2021), robustness evaluation of TDA features (Perez & Reinauer, 2022), authorship attribution of Japanese texts (Sakurai et al., 2025), out-of-distribution detection (OOD) (Pollano et al., 2024; Perez & Reinauer, 2022), and vulnerability detection in code (Snopov & Golubinskiy, 2024). Additionally, this framework is applied to the sentiment & semantic analysis task, specifically on human linguistic competence (i.e., grammatical acceptability judgment)  (Cherniavskii et al., 2022; Proskurina et al., 2023; Perez & Reinauer, 2022), dialog term extraction (Vukovic et al., 2022), and document coherence (Jain et al., 2024). Similarly, with the same framework, we observe a speech & music processing application (Tulchinskii et al., 2023). Finally, researchers attempt the model interpretation & analysis task, where Kostenok et al. (2023) uses the topological features extracted from the attention weights to estimate uncertainty in the encoder models, and Proskura & Zaytsev (2024) performs dynamic weighting for building ensemble models.
@@
Insight (Transformer-Attention). Given the wealth of information contained in the attention weights, researchers have applied topological features extracted from these weights to several tasks. The most interesting applications are observed on the sentiment and semantic tasks - grammatical acceptability judgment and document coherence analysis. Finally, we observe that for all the Transformer-based weights, the most explored subtask is deepfake text detection.
@@
ELMo Embeddings. ELMo embeddings are a type of word representation that captures both the meaning of words and their usage in context (Peters et al., 2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) performs text classification by applying this embedding for detecting fraudulent papers by examining their titles and abstracts. This is done in comparison of other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
@@
Insight (ELMo). ELMo embeddings is applied to one task - text classification by one researcher in comparison to other embeddings to determine the best numerical representation for extracting rich TDA features.
@@ 5.3.4 Symbolic Representations
Symbolic representations in the context of AI and cognitive science refer to the use of symbols such as letters, numbers, tokens, or abstract entities to represent concepts, objects, relationships, and rules within a system. These symbols can be manipulated according to predefined rules to perform reasoning, problem-solving, and decision-making. Symbolic representation contrasts with sub-symbolic representations, such as neural network-based embeddings, which do not explicitly use symbols or rules. This section then discusses the creation of symbolic representations by using principles of letter coding (PLC) and principles of speech sound coding (PSSC) which topological features are then extracted from.
PLC refers to rules and methods used to encode letters that fuel various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of Ukrainian tongue twisters (Yurchuk & Gurnik, 2023; Kovaliuk et al., 2024). These applications attempt two tasks, sentiment & semantic analysis and speech processing, respectively.
Yurchuk & Gurnik (2023) uses the PLC to create word embeddings for Ukrainian tongue twisters and extract topological features from such embeddings with persistent homology. This is to distinguish tongue twister from a simple narrative sentence using SVM and decision tree classifiers. Similarly, Kovaliuk et al. (2024) use PSSC for classifying spoken Ukrainian tongue twisters.
Insight (Symbolic). Symbolic representations are a novel technique for numerically representing text. The only applications of this technique are on Uranian tongue twisters, both for the sentiment & semantic analysis and speech & music processing tasks. This suggests that such a technique could be applied to processing low-resource languages, where current popular numerical representation may be insufficient.
@@ 5.3.5 Multi-Modal Representations
TDA features have also been extracted from other representations of NLP-related features, including multimedia data such as audio and video. In this section, the only task performed by researchers is speech & music processing, where applications include - studying human vowels and infant vocalizations (Bonafos et al., 2023; 2024), emotion recognition from audio speech (Gonzalez-Diaz et al., 2019) and audio in videos (Paluzo-Hidalgo et al., 2022), depression detection from audio clips (Tlachac et al., 2020), recognizing voiced and voiceless consonants in speech (Zhu et al., 2024), music classification (Bergomi, 2015; Sassone et al., 2022), and assessing the adversarial robustness of image-text multi-modal models by measuring topological consistency (Vu et al., 2025).
Insight (Multi-Modal). Multi-modal representations are another novel technique for numerically representing. The only task performed by researchers using this embedding is the speech & music processing task. This is because of data modality - audio data. All researchers show that topological features extracted from such embeddings experience high gains. Finally, the most interesting applications are the emotion recognition from audio speech and depression detection from audio clips.
@@
6 Case Studies of TDA for NLP Tasks
This section showcases the utility of TDA in non-trivial NLP tasks by applying persistent homology and mapper using both Theoretical (Section 6.1) and Non-theoretical (Section 6.2) approaches.
6.1 Theoretical Approaches
(Semantic) Topological Space.
One of the many applications of TDA in NLP includes projecting the semantic knowledge from texts into a topological space, creating a semantic topological space. While, this framework has yielded interesting findings as seen in Section 4.1.1, we are still unable to intuitively explain how and which semantic features TDA captures. To mitigate this, we aim to answer the question - What semantic features does TDA capture? Additionally, by attempting to answer this question, we illustrate a real-world application of TDA in NLP. We perform this analysis on a relevant topic - distinguishing between watermarked and non-watermarked LLM-generated texts.
We extract topological and linguistic features for a correlation test. This means that if the features are correlated, it could suggest that these TDA features extracted from text capture semantic features. For the analysis, we use a subset of the C4 dataset that has watermarked and non-watermarked LLM-generated texts [^5] https://huggingface.co/datasets/acmc/watermarked_c4_dataset. Next, we use the ripserripser python package to extract the topological features with persistent homology and textdescriptivestextdescriptives Python package [^6] https://github.com/HLasse/TextDescriptives?tab=readme-ov-file to extract the linguistic features. The text is initially represented numerically with BERT’s attention embeddings, then as an undirected graph, similar to Kushnareva et al. (2021) to extract topological features. Next, using the undirected graph as input, we calculate topological features in the 0-dimension, such as the number of non-zero values, maximum value, mean of values, number of values, and persistent entropy. Also, for the linguistic features, we extract entropy, perplexity, position of noun & adjective, coherence, flesh reading ease score & grade, number of stopwords, length of text, and number of unique words. Finally, we use distance correlation to measure the statistical dependence of the two non-linear sets of features. See Figures 8(a) & 8(b) for the correlation matrix heatmap. Many of the TDA features correlate with the linguistic features, with the highest correlation being the length of text (doc_length) and number of topological values in $\beta_{0}$ (betti_curve_0). Lastly, these results suggest that topological features can capture semantic features, however, a more comprehensive study is needed to make this a strong finding. In addition, such a finding can inform the non-theoretical approaches we adopt for real-world NLP tasks.
Topological “Shape” of Words
Using the same method above to represent the Watermarked vs. Non-Watermarked texts, we then visualize the “shape” of their data using Mapper to observe the similarity or distinctness between the text types. Using the same C4 dataset, we create Mapper diagrams using the $Mapper$ Python package [^7] https://github.com/scikit-tda/kepler-mapper/tree/master. Figures 9(a) and 9(b) depict the Mapper plots of the Watermarked and Non-Watermarked texts, respectively. Represented as intuitive structured graphs, the figures show that texts across the two categories are topologically similar, although the watermarked text has slightly more nodes and edges than the non-watermarked texts, suggesting more connectivity. Thus, we can conclude that these texts are topologically similar, which could suggest that the watermarking technique does not significantly deviate from the style and quality of the non-watermarked version. In other words, such findings could indicate that the watermarking technique does not break the semantics, style, and syntax of the non-watermarked pre-trained model. However, to further explain these results, a more comprehensive study is required to interpret topological features extracted from text linguistically. This further stresses the use of the theoretical study of TDA as a powerful analytical tool as described in Section 6.1.